A simple pendulum swings freely and rises at the end of its swing to a position 8.5 cm above its lowest point. What is its speed at its lowest point?


Answer using Grade 11 Conservation of Energy and Momentum please!



1
Expert's answer
2020-05-07T10:19:16-0400

Here total mechanical energy conservation will take place

"K_i+P_i=K_f+P_f"

"0+m\times10\times0.085=0.5mv^2+0\\v^2=0.85\\v=0.92 m/s"